OpenStudy (anonymous):

Please help I have to have a math assignment turned in today and I just can't figure this out! Help solve this equation 7b = 3 - 3b = 3 - 3(b+4). Thanks in advance.

OpenStudy (babystar16):

its going to be a decimal so round it

OpenStudy (babystar16):

b=1.714285714

OpenStudy (babystar16):

7b=3-3b=3-3b-12 7b=3-3b=-9-3b 7b=12-3b=-3b 7b=12 b=1.714285714

OpenStudy (babystar16):

thats how i got it

OpenStudy (anonymous):

thanks.

OpenStudy (babystar16):

ure welcome